### v1

# Scaffold-BPE: Enhancing Byte Pair Encoding with Simple and Effective Scaffold Token Removal

Abstract
Byte Pair Encoding (BPE) serves as a foundation method for text tokenization in the Natural Language Processing (NLP) field. Despite its wide adoption, the original BPE algorithm harbors an inherent flaw: it inadvertently introduces a frequency imbalance for tokens in the text corpus. Since BPE iteratively merges the most frequent token pair in the text corpus while keeping all tokens that have been merged in the vocabulary, it unavoidably holds tokens that primarily represent subwords of complete words and appear infrequently on their own in the text corpus. We term such tokens as Scaffold Tokens. Due to their infrequent appearance in the text corpus, Scaffold Tokens pose a learning imbalance issue for language models. To address that issue, we propose Scaffold-BPE, which incorporates a dynamic scaffold token removal mechanism by parameter-free, computation-light, and easy-to-implement modifications to the original BPE. This novel approach ensures the exclusion of low-frequency Scaffold Tokens from the token representations for the given texts, thereby mitigating the issue of frequency imbalance and facilitating model training. On extensive experiments across language modeling tasks and machine translation tasks, Scaffold-BPE consistently outperforms the original BPE, well demonstrating its effectiveness and superiority.
[…]
284
1 Introduction
In recent years, large language models (LLM) [33, 8, 42] have made a substantial impact on Natural Language Processing (NLP). Those models, which are usually extremely huge in parameter scales [42, 34, 9], demand managing and processing vast vocabularies efficiently during both training and inference stages, and pose great challenges to the communities. To address that issue, Byte Pair Encoding (BPE) was applied to construct vocabularies. BPE was initially introduced as a data compression algorithm [14]. It iteratively merges the most frequent pairs of bytes or characters in a dataset until a desired vocabulary size is reached. The capability of BPE to break down words into more manageable subword units allows for more flexible and semantically complete representations of input data. Therefore, the BPE technique avoids the out-of-vocabulary problem [40], a prevalent issue where unknown words disrupt the functioning of NLP models. Consequently, BPE has been attracting much attention across the community [31, 49, 38]. In the NLP field, BPE is a cornerstone for numerous applications, like machine translation [40, 32, 49, 17], language understanding [27, 19], and even Large Language Model (LLM) training [33, 8, 5, 42, 50, 34, 52].
Figure 1: Two types of tokens trained by the original BPE on the Pile dataset: one type, such as “age”, can appear frequently by itself, while the other type, such as “zona”, mostly appears as a subword of “Arizona” and thus has a low individual appearance frequency. The value on the horizontal axis represents the percentage of the frequency of the token relative to the total frequency of all tokens containing the specific token. For “age”, we only visualize the top tokens.
Since its inception, BPE has undergone various modifications to better suit the needs of complex natural language processing tasks. Existing works [45, 21, 12, 49, 17, 36, 39, 18, 32, 20, 38] have thoroughly investigated to improve BPE from multiple perspectives. A significant trend of research has focused on identifying the optimal vocabulary size that BPE should target [49, 17, 36, 39, 18, 32, 12]. These investigations reveal that a meticulously calibrated vocabulary size can reduce the computational load while improving the linguistic accuracy of the models. Besides, several works [32, 20, 38] have looked into the encoding process of BPE. By optimizing the encoding paths of tokens, these methods capture a wider range of linguistic phenomena and thus improve model performance across diverse datasets.
However, existing studies have overlooked a significant limitation inherent in the BPE method: the iterative merging process can lead to an imbalance in token frequencies by including low-frequency subwords as tokens in vocabulary. For example, as illustrated in Figure 1, in the commonly used Pile dataset [15] for training LLMs [5, 48, 41] tokenized by the original BPE, the token “zona” mostly appears as a subword within the token “Arizona” rather than as an independent, high-frequency token. Despite its lower standalone frequency, BPE includes “zona” in the final vocabulary because it is the “intermediate token” to derive the frequent token “Arizona”. We define such intermediate tokens that are crucial for constructing longer frequent tokens but do not appear frequently on their own as Scaffold Tokens. As we observe, a 32K vocabulary (size applied by LLaMA series [42, 43]), trained on the Pile dataset [15] with the original BPE [40], contains about 6.07% of scaffold tokens. As depicted in Figure 2, a natural frequency imbalance arises between these scaffold tokens and actual high-frequency tokens. Prior studies [26, 41] have highlighted that such disparities in token frequencies can result in imbalanced learning difficulties across different tokens. Scaffold tokens, due to their lower individual appearance frequencies, are notably harder to learn for models. To address that issue, we propose enhancements to the BPE algorithm aimed at mitigating the frequency imbalance and ensuring a more equitable learning process for all tokens.
Figure 2: Sorted token frequencies in descending order of the original BPE and Scaffold-BPE.
Specifically, we propose the simple and effective Scaffold-BPE with a dynamic scaffold tokens removal mechanism, which is parameter-free, computation-light, easy-to-implement, and widely effective. Generally, the proposed Scaffold-BPE expands the original BPE vocabulary. In the training stage, Scaffold-BPE dynamically marks tokens with lower individual appearance frequencies as scaffold tokens in each iteration. Note that the scaffold tokens are not actual tokens in the vocabulary and do not appear in the tokenized sequences after encoding. In the encoding stage, the Scaffold-BPE firstly utilizes all tokens in the expanded vocabulary to generate the token representations for the given texts, which is termed as a Scaffolding process. Then, the Scaffold-BPE ensures the absence of all scaffold tokens in the token representation by demolishing them into their shortest non-scaffold-tokens sequence, which is termed as a Demolishing process. Thanks to such modifications, Scaffold-BPE can remove scaffold tokens from the final token representations fed into models for more balanced token occurrences, thus leading to more sufficient learning and better performance of models.
We conduct extensive experiments on language modeling tasks and machine translation tasks. First, experiments on six widely used language modeling benchmarks including HellaSwag [51], OpenBookQA [29], PIQA [6], SIQA [37], StoryCloze [30], and Winogrande [35] demonstrate that Scaffold-BPE consistently outperforms the original BPE. Besides, experiments on WMT’14 English-German and WMT’14 English-France machine translation tasks [7] also indicate that Scaffold-BPE outperforms the original BPE.
Overall, our contributions are three-fold:
  • We observe that the iterative training process of BPE incorporates tokens with imbalanced frequency distributions into the vocabulary, which we term scaffold tokens.
  • We propose Scaffold-BPE, which can remove scaffold tokens from the final token representations by dynamically marking scaffold tokens in the training process and temporarily utilizing scaffold tokens in the encoding process. Scaffold-BPE is parameter-free, computation-light, easy-to-implement, and widely effective, preserving the simplicity and clarity of BPE.
  • Extensive experiments demonstrate that Scaffold-BPE surpasses the original BPE on language modeling and machine translation tasks, proving its effectiveness and robustness in the NLP field.
2 Related Work
2.1 Byte Pair Encoding
In the early stages of neural model development, researchers primarily constructed models based on word-level vocabularies [11, 44, 53], which showed considerable success.
However, those models often struggled with the challenge of rare word handling due to the inherent limitations of word-level vocabulary size. In response, the academic community has explored numerous innovative strategies for vocabulary development, including methods based on bytes [46], characters [11, 25, 1], and subwords [40, 24].
Among those, Byte Pair Encoding (BPE) [40] stands out for its effective creation of subword vocabulary. Its design philosophy is notably straightforward. During the training process, the corpus is initially split into a sequence of the smallest unit tokens (i.e., character tokens [40] or byte tokens [46]). The algorithm iteratively finds the most frequent token pairs in the sequence, merges them into a new token, and adds it to the vocabulary until it reaches a predetermined size. The vocabulary is then utilized during the encoding phase to represent any text. It reduces token sparsity and enhances feature identification in related words sharing an identical subword, without losing rare words. Recent advancements like BPE-dropout [32] and optimal vocabulary size search [49, 17, 39, 36] continue to enrich BPE development in neural models.
However, previous works did not take into account a fundamental flaw of BPE: during the training process of BPE, existing tokens are merged into new tokens, resulting in a decrease of their individual frequencies that are not covered by the new tokens. That leads to the inclusion of some lower-frequency tokens in the vocabulary, hindering the inclusion of other actual high-frequency tokens, thus resulting in an imbalance of token frequencies and wastage of the vocabulary. To address that issue, this paper introduces Scaffold-BPE, which has a dynamic scaffold token removal mechanism that ensures the tokens fed into models are actual high-frequency tokens.
2.2 Language Models
Figure 3: Illustration of one iteration in the Scaffold-BPE training process.
Language models are designed to predict the probability distribution of a token sequence. Following GPT-3 [8], which features 175 billion parameters and demonstrates versatility across a wide range of applications, there has been a significant push towards developing large generative language models like Gopher [34], PaLM [9], GaLM [13], OPT [52], and LLaMA [42]. Such a surge in development has greatly advanced the fields of natural language understanding and generation.
However, as a mainstream tokenizer used by language models [33, 8, 5, 42, 50, 34, 52], BPE has inherent flaws of imbalanced token frequencies, which can impact the training of language models. First, the low-frequency tokens are updated infrequently during the training process, leading to poorer performance of those tokens. Second, the presence of low-frequency tokens in the vocabulary prevents other actual high-frequency tokens from being included. Together, the two points above further impact the performance of language models on downstream tasks. This paper demonstrates that by using our Scaffold-BPE algorithm as the tokenizer, language models can achieve a consistent improvement on downstream tasks.
3 Methodology
To enhance the performance of the original BPE, we propose Scaffold-BPE to remove the scaffold tokens introduced by the original BPE. Our Scaffold-BPE is simple and straightforward. In the training process, the Scaffold-BPE dynamically marks scaffold tokens in the vocabulary at each iteration, and finally yields an expanded vocabulary consisting of both normal tokens with the amount equaling the predetermined vocabulary size and several scaffold tokens. In the encoding process, apart from using the normal tokens, Scaffold-BPE temporarily uses previously marked scaffold tokens as intermediate tokens to merge into longer normal tokens.
3.1 Training Process
Algorithm 1 Scaffold-BPE Training Algorithm
Text Corpus $C$, Desired Vocabulary Size $N$
Initialize an expanded vocabulary $E$, consisting of a normal-token vocabulary $V$ and a scaffold-token vocabulary $S$
Split $C$ into a list of characters/bytes (denoted as $L$)
Initialize a priority queue $Q$(token pair, $f(\cdot)\downarrow$) within $L$
while $|V|$ < $N$ do
     $(a,b)$leftarrow pop $Q_{head}$
     Merge pair $(a,b)$ into a new token $t$
     if $t$ in $S$ then $\triangleright$ $t$ may be a previously marked scaffold token
         Scaffold$(t)\leftarrow$ False
         continue
     end if
     Add $t$ to $E$ as a normal token
     Replace all of $(a,b)$ in $L$ with $t$
     Update $Q$
     /******* Scaffold-BPE Modification Begins *******/
     for each $p$ in {$a$, $b$do
         $f(p)\leftarrow f(p)-f((a,b))$
         if $p$ in $V$ and $f(p)$ < $f(Q_{head})$ then
              Scaffold$(t)\leftarrow$ True
              Push $p$ back to $Q$
         end if
     end for
     /******** Scaffold-BPE Modification Ends ********/
end while
return $E$ (with $V$ and $S$ both included)
The original BPE is trained on a text corpus $C$ with a predefined vocabulary size $N$. After training, BPE returns a vocabulary $V$ consisting of $N$ tokens. For simplicity, $C$ is firstly split into a sequence of smallest unit tokens (denoted as $L$), with each token being a single character/byte. We define $a$, $b$ as two tokens, $(a,b)$ as a token pair, and $f(\cdot)$ as the frequency of a token or token pair within $L$. BPE is trained iteratively. In each iteration, BPE identifies the token pair with the highest frequency:
$(a,b)=\arg\max_{(x,y)\in L}f((x,y))$ (1)
BPE then merges (i.e., concatenates) them into a new token $t$, and includes $t$ in $V$. Then BPE updates $L$ via replacing all $(a,b)$ with $t$, and restarts the process again. The iterative process of identifying the most frequent token pair $(a,b)$ can be accelerated using a priority queue $Q$. At the beginning of the training process, all token pairs in $L$ are pushed into $Q$ with a descending order of frequency. And after the token pair $(a,b)$ is merged into $t$ in each iteration, BPE updates the frequency and rank of token pairs related to all indexed occurrences of $(a,b)$. For instance, given $(a,b)$ in a context of $\ldots,u,a,b,v,\ldots$ in $L$, when $(a,b)$ is replaced with $t$, the frequency of $(u,a)$ or $(b,v)$ would decrease by 1, and meanwhile that of $(u,t)$ or $(t,v)$ would increase by 1. With the occurrences of all token pairs being indexed, there is no need to scan $L$ again and re-count the frequencies of all candidate token pairs for a new iteration. After updating the adjacent token pairs related to $(a,b)$ (i.e, $t$), the frequencies of token pairs like $(u,a)$ or $(b,v)$ would be updated in $Q$, and meanwhile the new candidate token pairs $(u,t)$ and $(t,v)$ would also be pushed into $Q$ with their corresponding frequencies.
The Scaffold-BPE expands the vocabulary $V$ to an expanded vocabulary $E$, and assigns an attribute (denoted as $\text{Scaffold}(\cdot)$) to each token in the vocabulary indicating whether it is a scaffold token or not. Thus, the expanded vocabulary $E$ comprises two types of tokens. We denote all the non-scaffold-tokens by $V$, which, as with the original BPE, are the tokens actually used in model training.
$V=\{t\in E\mid\text{Scaffold}(t)=False\}$ (2)
Additionally, we denote all the scaffold tokens by $S$, which are not fed into the model, nor do they appear in any token representations after encoding.
$S=\{t\in E\mid\text{Scaffold}(t)=True\}$ (3)
They only serve as intermediate tokens to aid in the training and encoding processes of Scaffold-BPE. Therefore, when calculating the size of the vocabulary, the count of scaffold tokens is not included; only the number of tokens in $V$ is considered.
Initially, a token pair is merged and added to $E$ due to its high frequency. Similarly, Scaffold-BPE marks a token as a scaffold token when its frequency decreases. Throughout the entire training process of BPE, $f(a)$ and $f(b)$ only decrease when the token pair $(a,b)$ is merged into a new token $t$. Therefore, as presented in Algorithm 1, Scaffold-BPE introduces an additional step at the end of each iteration, utilizing the reduced $f(a)$ and $f(b)$ to evaluate whether $a$ and $b$ remain high-frequency. If they are no longer considered high-frequency, they would be marked as scaffold tokens. Naturally, the token pair at the head of the priority queue $Q$ (denoted as $Q_{head}$) is the next candidate to be added to the vocabulary. Then $f(Q_{head})$ is a natural frequency delimiter between in-vocabulary and out-vocabulary tokens. Therefore, if $f(a)$ (or $f(b)$) $<f(Q_{head})$, $a$ (or $b$) is marked as a scaffold token, which means it is not included by $V$. Notably, such an additional step leverages the inherent mechanism of BPE without introducing any additional hyper-parameters, maintaining the simplicity and clarity of BPE. Moreover, $f(Q_{head})$ is dynamically adjusted in each iteration, ensuring that Scaffold-BPE can adaptively identify scaffold tokens at any iteration step. Furthermore, scaffold tokens are not permanently marked. They are pushed back into $Q$, reserving the possibility of being ranked top at the priority queue and re-integrated into $V$ in a future iteration.
3.2 Encoding Process
The encoding process of the original BPE encodes a text $T$ into a token representation (i.e., $R$) using the vocabulary $V$ generated by BPE training. Firstly, $R$ is a sequence of smallest unit tokens (i.e., character/byte tokens), obtained by splitting $T$. And then, following the order of tokens in $V$ as merging priority (i.e., tokens added earlier have higher frequency and thus are assigned higher priority to be merged into), token pairs in $R$ are iteratively merged to build the final representation.
Algorithm 2 Scaffold-BPE Encoding Algorithm
A Text $T$, Expanded Vocabulary $E$
Split $T$ into a character/byte token representation (denoted as $R$)
while True do
     /**************** Scaffolding Begins ****************/
     Identify all possible merges $M$ using $E$, ignoring token types
     /***************** Scaffolding Ends *****************/
     if $M$ is empty then
         break
     end if
     Select $m$ which is ranked before the others in $E$ from $M$
     apply $m$ to $R$
end while
/***************** Demolishing Begins *****************/
Demolish all scaffold tokens in $R$ into its shortest non-scaffold child token sequence
/****************** Demolishing Ends ******************/
return $R$
Similarly, the modifications of Scaffold-BPE in the encoding process are very simple. Compared to the original BPE, the expanded vocabulary $E$ is utilized. Both normal tokens and scaffold tokens are merged according to their rank in $E$. Consequently, during the encoding process, the count of different tokens used actually exceeds the predefined vocabulary size (i.e., $N$). And scaffold tokens are employed as intermediate tokens to merge into longer tokens. We term that mechanism as Scaffolding, as shown in Algorithm 2.
Figure 4: The encoding process (i.e., Scaffolding and Demolishing) for the word “zona" and “Arizona".
When no more token pairs can be merged in $R$, the original BPE returns $R$ as the final result. However, due to the introduction of the Scaffolding mechanism in Scaffold-BPE, $R$ may contain scaffold tokens from $S$, potentially increasing the variety of tokens beyond the predefined vocabulary size and exceeding the range of word embeddings that the model can map. To address it, Scaffold-BPE adds one additional step termed as Demolishing at the end of the encoding process. Scaffold-BPE demolishes all scaffold tokens in $R$ into their shortest non-scaffold child token sequences, ensuring that $R$ only consists of tokens from $V$. For example, as shown in Figure 4, the remaining “zona” in $R$ is demolished into “zon” and “a”. After the Demolishing step, Scaffold-BPE returns the final token sequence representation (i.e., $R$) for $T$. Since the shortest non-scaffold child token sequences for all scaffold tokens can be precomputed and stored during the training process, the time complexity of demolishing one token is $O(1)$, making its impact on encoding efficiency negligible.
4 Experiments
We employ the recently well-attended language modeling tasks to validate the effectiveness of the Scaffold-BPE.
4.1 Experimental Setup
Datasets. Our models are trained on the Pile [15] dataset, an 825.18 GiB English text dataset designed for training large scale language models. The Pile is composed of 22 diverse and high-quality datasets, the models trained on which significantly outperform both raw and filtered Common Crawl [10] models. The data distribution for our model training is identical to those described in the original work [15].
Tokenizer. We train two 32K vocabularies (size applied by LLaMA series [42, 43]) using the original BPE and Scaffold-BPE, respectively. The training text corpus is sampled from the Pile dataset with an identical data distribution. Following GPT-2 [33], we pre-segment the text using its regular expression.
Table 1: Model sizes and architectures.
model size dimension $n$ heads $n$ layers batch size seq length
468M 1024 16 24 1024 1024
1.2B 2048 8 16 2048 1024
6.7B 4096 32 32 2048 2048
Model. We train three generative language models with 468M, 1.2B, and 6.7B parameters, respectively. Specifically, the architectures of the 468M-parameter and the 1.2B-parameter models, including the dimensionality of hidden states, the number of layers, etc., are identical to those of the 410M-parameter and the 1.0B-parameter models outlined in Pythia [5]. The minor differences in parameter sizes are attributed to the variations in vocabulary size in the embedding layer. As for the 6.7B-parameter model, its architecture is identical to LLaMA-7B [42]. The corresponding hyperparameters for each model can be found in Table 1.
Training. Following LLaMA [42], we use the AdamW optimizer [28] with a learning rate of $3.0\times 10^{-4}$, $2k$ warmup steps, and a cosine learning rate decay schedule. Following the pretraining settings of previous works [48, 41] and limited by our computation budget, by default all models are pretrained with 100B tokens. Note that the volume of corresponding text data contained in an equal amount of tokens is slightly different between the two tokenizers. Considering model training efficiency and commonly used criteria (i.e., the token amount) of computation budget in LLM training, we still compare experiments in the setting of an equal amount of training tokens. In Section 5.3, we further analyze both tokenizers in the setting of an equal amount of training text volume.
Evaluation. For fair comparisons, we utilize the open-source pipeline lm-evaluation-harness [16] for evaluation [5, 41].
4.2 Experimental Results
Table 2: At varying model scales, the zero-shot and few-shot performance (i.e., accuracy %) on common sense reasoning benchmarks.
HellaSwag OpenBookQA PIQA SIQA StoryCloze Winogrande
468M
0-shot Original BPE 40.81 30.20 66.74 41.62 63.70 53.12
Scaffold-BPE 41.23 30.64 66.94 42.14 63.56 53.20
5-shot Original BPE 40.99 29.20 66.60 43.88 63.00 53.40
Scaffold-BPE 41.19 29.88 66.79 44.98 63.52 53.48
1.2B
0-shot Original BPE 47.21 30.68 69.25 41.77 65.92 55.60
Scaffold-BPE 47.50 31.12 69.52 41.97 66.31 55.53
5-shot Original BPE 47.39 32.36 69.61 46.03 65.79 55.33
Scaffold-BPE 47.64 32.40 69.84 46.58 66.57 56.15
6.7B
0-shot Original BPE 60.84 36.56 73.72 43.00 71.20 61.50
Scaffold-BPE 60.50 37.12 73.99 43.91 71.61 61.97
5-shot Original BPE 61.24 36.24 74.40 49.58 72.63 62.83
Scaffold-BPE 61.34 38.08 74.44 49.89 72.76 63.47
Common Sense Reasoning. Our analysis incorporates six benchmark datasets recognized for evaluating common sense reasoning including HellaSwag [51], OpenBookQA [29], PIQA [6], SIQA [37], StoryCloze [30], and Winogrande [35]. We present the performance of our model, focusing on accuracy in both zero-shot and few-shot scenarios.
As shown in Table 2, we can observe that the Scaffold-BPE consistently outperforms the original BPE on different setups with different model sizes. Specifically, for all models on 5-shot settings, Scaffold-BPE consistently achieves superior performance compared to the original BPE. As for the 0-shot settings, Scaffold-BPE still yields higher performance on 5 out of the 6 datasets.
Such results clearly demonstrate that although the modifications are simple, our proposed Scaffold-BPE is convincingly effective. We attribute it to that Scaffold-BPE can encode text into tokens with a more balanced frequency distribution, which can help language models to learn all tokens more thoroughly.
Table 3: Zero-shot and few-shot exact match performance of 6.7B-parameter language models on closed-book question-answering benchmarks.
TriviaQA WebQuestions
0-shot Original BPE 11.32 3.84
Scaffold-BPE 12.90 5.11
5-shot Original BPE 23.50 16.44
Scaffold-BPE 23.68 17.13
Closed Book Question Answering. For the task of closed book question answering [8, 42, 41], we evaluate the performance of the largest 6.7B-parameter models with different tokenizers on two benchmark datasets, i.e., TriviaQA [22] and WebQuestions [4]. We report the exact match performance for the zero-shot and few-shot settings in Table 3. It can be seen that language models trained with the proposed Scaffold-BPE achieve superior performance in both settings, which demonstrates that Scaffold-BPE can enhance model performance across different types of downstream tasks.
5 Discussion
5.1 Robustness In Language Modeling Tasks
Various Vocabulary Size. Depending on the size of the training corpus, the diversity of the languages, the size of the model, and the types of tasks, different vocabulary sizes are set in practice. Therefore, to validate the robustness of Scaffold-BPE across various vocabulary sizes, in addition to the 32K vocabulary [42, 34], we also trained two vocabularies sized at 64K [3, 2] and 128K [50]. The experimental setup is identical to that of the 468M-parameter model mentioned in Section 1.
Figure 5: The average frequencies of the scaffold tokens and the new actual high-frequency tokens that replace the removed scaffold tokens in vocabularies of 32K, 64K and 128K.
As shown in Figure 5, Scaffold-BPE can replace scaffold tokens in vocabularies with actual high-frequency tokens, which significantly increases the average frequencies of those tokens. The frequency improvements are 76.40%, 68.58%, and 58.99% for the 32K, 64K, and 128K vocabulary sizes, respectively. The enhancement in token frequency distribution effectively promotes the learning of those tokens, which can contribute to better model performance across various tasks.
As shown in Table 4, the experiments demonstrate that Scaffold-BPE consistently outperforms the original BPE across all vocabulary sizes, which indicates that Scaffold-BPE is not sensitive to vocabulary size. Its algorithmic design enables it to adaptively identify and remove scaffold tokens across any vocabulary size, without the need for manually designed or heavily-tuned hyperparameters.
Table 4: At varying vocabulary sizes, the zero-shot and few-shot performance (i.e., accuracy %) on common sense reasoning benchmarks.
HellaSwag OpenBookQA PIQA SIQA StoryCloze Winogrande
64K
0-shot Original BPE 41.61 30.00 67.70 41.46 63.45 52.77
Scaffold-BPE 42.09 31.20 67.71 41.80 63.83 52.80
5-shot Original BPE 42.05 30.08 67.58 44.22 63.98 53.25
Scaffold-BPE 42.64 29.92 67.70 45.16 64.40 53.67
128K
0-shot Original BPE 43.59 31.60 67.97 41.72 65.07 53.43
Scaffold-BPE 43.41 31.92 68.80 42.18 65.23 54.30
5-shot Original BPE 43.40 29.24 68.12 45.04 65.08 53.53
Scaffold-BPE 43.45 31.36 69.16 45.91 65.59 54.90
More Training Tokens. According to the Scaling Law, the loss scales as a power-law with model size, dataset size, and the amount of compute used for training [23]. To demonstrate the effectiveness of our Scaffold-BPE with more training tokens, we continue training the 468M-parameter models up to 300B tokens [52, 5, 34, 8].
As shown in Table 5, the experiments demonstrate that Scaffold-BPE consistently outperforms the original BPE at 300B training tokens. The result well indicates that in the era of increasingly large training datasets for Large Language Models, our Scaffold-BPE can effectively enhance the capabilities of those models through simple modifications to the original BPE.
Table 5: At 300B training tokens, the zero-shot and few-shot performance (i.e., accuracy %) on common sense reasoning benchmarks.
HellaSwag OpenBookQA PIQA SIQA StoryCloze Winogrande
0-shot Original BPE 45.56 30.40 68.35 42.08 65.25 54.40
Scaffold-BPE 45.69 30.60 69.53 42.37 65.36 54.95
5-shot Original BPE 45.72 30.72 68.31 46.10 65.58 54.96
Scaffold-BPE 45.78 31.04 69.53 46.15 65.95 56.13
5.2 Task Insensitive, Language Insensitive And Architecture Insensitive
Although the development of large language models is burgeoning and they are increasingly applied across various scenarios, many applications still prefer using conventional models due to their lower training and inference costs. In the field of Natural Language Processing, BPE was initially applied to machine translation tasks [40], which typically present an open vocabulary challenge and involve substantial textual variation between two languages. Consequently, numerous improvements to BPE have been extensively validated on machine translation tasks [32, 49, 20, 45, 12, 46, 36].
Therefore, to validate the versatility of the Scaffold-BPE method, we additionally conduct evaluations on machine translation tasks. We replicate the experimental setup of the prior work [31] which uses 32K vocabularies for the WMT’14 English-German dataset and 40K vocabularies for the WMT’14 English-French dataset [7]. For fair comparisons, We do not pre-segment the text using regular expressions. We train the “big" transformer models [44, 31] to convergence and average model parameters from the last 10 checkpoints [31].
Table 6: BLEU on newstest2014 for WMT English-German (En–De) and English-French (En–Fr).
En-De En-Fr
Original BPE [31] 29.3 43.2
Scaffold-BPE 29.8 43.8
As shown in Table 6, Scaffold-BPE outperforms the original BPE in machine translation tasks, which demonstrates that Scaffold-BPE is not specific to language modeling tasks and can be applied to a wider range of tasks like language understanding, summarization and text classification.
Besides, experiments conducted with English-German and English-French language pairs demonstrate that Scaffold-BPE is language insensitive. Scaffold-BPE is capable of identifying and removing the scaffold tokens introduced by the original BPE across different languages.
Furthermore, prior experiments on language modeling tasks are carried out on decoder-only architecture. For the machine translation tasks, we utilize the classic encoder-decoder architecture [44]. The exceptional performance of Scaffold-BPE confirms its architecture insensitivity, indicating its applicability across a wider range of neural network architectures.
5.3 Better Text Representation
Higher Entropy, Lower Redundancy. Scaffold-BPE can alleviate the imbalance in token frequency, which can lead to an increase in information entropy. We measure Shannon Entropy and Redundancy [18] over token representations of texts obtained with the original BPE and our Scaffold-BPE. Both take as input a text $T$ with a vocabulary of (normal) tokens $V=\{t_{1},t_{2},...,t_{V}\}$ of size $|V|$.
Entropy $H$ is a measure of the average information. Where the probability of a token $p(t)$ is estimated using the so-called maximum likelihood method (i.e., its relative frequency in the text). Higher values of Entropy indicate higher complexity (less predictability).
$H(T)=-\sum_{i=1}^{V}p(t_{i})\log_{2}p(t_{i})$ (4)
The Redundancy $R$ quantifies how close the empirically estimated entropy is to the maximum value it can take.
$R(T)=1-\frac{H(T)}{max\{H(T)\}}=1-\frac{H(T)}{\log_{2}|V|}$ (5)
As shown in Table 7, taking the 32K vocabulary as an example, our Scaffold-BPE can encode Pile dataset [15] with higher Entropy and lower Redundancy. Consequently, tokens in the vocabulary of our Scaffold-BPE have more balanced appearing probabilities. According to Su et al. [41], our vocabulary with balanced token appearances mitigates the learning imbalance problem, resulting in more sufficient learning towards the text corpus, thus achieving better performance.
Table 7: Entropy and Redundancy on tokenized Pile dataset.
[…]
Original BPE 11.2382 0.2491
Scaffold-BPE 11.2443 0.2487
Table 8: Compression Rate (bytes per token) on the Pile dataset and the WMT dataset.
Pile-32K Pile-64K Pile-128K En-De En-Fr
Original BPE 3.879 4.083 4.228 4.830 5.012
Scaffold-BPE 3.889 4.089 4.232 4.861 5.042
Table 9: In the condition of exactly 388 GiB training text, the zero-shot and few-shot performance (i.e., accuracy %) on common sense reasoning benchmarks.
HellaSwag OpenBookQA PIQA SIQA StoryCloze Winogrande
0-shot Original BPE 40.80 30.21 66.74 41.68 63.65 53.12
Scaffold-BPE 41.24 30.66 66.92 42.15 63.55 53.22
5-shot Original BPE 40.97 29.22 66.63 43.86 63.03 53.44
Scaffold-BPE 41.20 29.87 66.78 45.01 63.54 53.50
Higher Compression Rate. Besides the performance of models on tasks, the compression rate is a metric to measure the effectiveness of a tokenizer. A higher compression rate means that fewer tokens are required to represent the same corpus. As shown in Table 8, Scaffold-BPE, utilizing a dynamic scaffold tokens removal mechanism, retains more actual high-frequency tokens in the final vocabulary. Therefore it can achieve a higher compression rate on the corpus.
Besides, considering model training efficiency and commonly used criteria (i.e., the token amount) of computation budget in LLM training, experiments in Section 4.2 and 5.1 are compared in the setting of an equal amount of training tokens. To eliminate the impact of different amounts of training text caused by different compression rates on experiment results, we additionally train two 468M-parameter models on exactly 388 GiB training text ($\approx$ 100B tokens). As shown in Table 9, the Scaffold-BPE consistently outperforms the original BPE, demonstrating that the effectiveness of Scaffold-BPE is not merely obtained by allowing models to digest more data in the same computation budget. Our Scaffold-BPE addresses the issue of token frequency imbalance, allowing models to learn the information contained in tokens more sufficiently, thus achieving better performance.
Better Uniformity of Learned Embeddings. Prior works have analyzed the embedding space learned by a model [32] and found that better uniformity prefers a token embedding space that preserves maximal information [47]. To demonstrate our Scaffold-BPE can mitigate token frequency distribution imbalance, thus leading to a better learned token embedding space with better uniformity, we visualize the token embeddings in the 6.7B-parameter models, following Provilkov et al. [32]. As shown in Figure 6, the embeddings of scaffold tokens learned via the original BPE are more clustered, which means they are not well learned. On the contrary, the embeddings of new tokens introduced by Scaffold-BPE after removing scaffold tokens have better uniformity, which are more evenly distributed across the semantic space. Therefore, models trained with Scaffold-BPE can achieve better performance.
(a) The Original BPE
(b) The Scaffold-BPE
Figure 6: Visualization of the learned embeddings of tokens in the respective vocabulary generated by the original BPE and Scaffold-BPE.
6 Conclusions
In this paper, we present our observation of tokens with imbalanced frequencies in BPE vocabulary, which we term scaffold tokens. Those scaffold tokens, while integral to the formation of longer tokens, do not represent actual frequent tokens in the corpus and affect NLP task performance. To address that issue, we propose Scaffold-BPE, which can remove scaffold tokens from the final token representations by dynamically marking scaffold tokens in the training process and temporarily utilizing scaffold tokens in the encoding process. The Scaffold-BPE is parameter-free, computation-light, easy-to-implement, and widely effective, well preserving the simplicity and clarity of BPE. Through extensive experiments, including varying model sizes, varying vocabulary sizes and extending training tokens, Scaffold-BPE demonstrates its robustness and superiority over the original BPE across a variety of natural language processing tasks. Our work underscores the importance of continual refinement in tokenization methods for improving the overall efficiency and effectiveness of models in the natural language processing field.
7 Limitations
In the proposed Scaffold-BPE, the modifications to the training and encoding of the original BPE are simple and straightforward. Therefore Scaffold-BPE may be combined with other enhancements such as optimal vocabulary size search and novel encoding methods to achieve further improvements. We leave the investigation to our future research.